OpenStudy (anonymous):

what is the cube root of 3456

OpenStudy (tkhunny):

You can always guess. 14*14*14 = 2744 15*15*15 = 3375 16*16*16 = 4096 Well, that was pretty good guessing. 15*15*15 = 3375 <== From Above 15.1^3 = 3442.951 15.2^3 = 3511.808 15.1^3 = 3442.951 <== From Above 15.11^3 = 3449.795831 15.12^3 = 3456.649728 <== Just a hair over the integer portion. 15.119^3 = 3455.963930159 15.1191^3 = 3456.032505860871 15.11905^3 = 3455.998217896542625 15.119051^3 = 3455.998903653606689651 15.119052^3 = 3455.999589410761468608 <== 3 9's is pretty good. 15.1190525^3 = 3455.999932289372875953125 <== 4 9's! 15.1190526^3 = 3456.000000865097878851576 <== 5 0's! (Didn't count the last 0 because of rounding.) Of course, using the \(\left[x^{y}\right]\) button on your calculator might be faster. 3 4 5 6 \(\left[x^{y}\right]\) ( 1 / 3 ) = 15.119052598738477977206527287339 Wasn't that fun?